decide whether the pair of lines is parallel, perpendicular or neither
3x+5y=5
5x+3y=7

Answer by nabla(475)

Solve each for y.
1) y=-3x/5+1
2) y=-5x/3+7/3
So slopes (m for y=mx+b) are:
1) m=-3/5
2) m=-5/3
So these are reciprocals of eachother, however, for perpendicularity (or orthogonality ) we require negative reciprocals. So we say neither.
In fact, the graph shows a moderately compressed cross, when we would want a "perfect" cross, if you get my meaning.